# HelixFold-Multimer: Elevating Protein Complex Structure Prediction to New Heights

### v1

@@ 1.4 Sequence Identity
To comprehensively grasp the impact of sequence similarity on model performance, we employed sequence identity as the similarity metric. Sequence identity is defined as the percentage of residues in the evaluation set chain that match those in the training set chain. We utilized the default configuration settings of the MMSeqs2[17] software suite for this computation.
@@ 2.1 General Version
Heteromeric protein complexes released from the period between January 16th, 2022, and December 12th, 2022, were gathered from Protein Data Bank (PDB) [23] as the evaluation set. Proteins with more than 1400 residues are excluded. These test samples are not present in the training set. The evaluation set for heteromeric proteins consists of 194 complexes.
To assess the models on protein-peptide complexes, we derived a subset of protein complexes from the heteromeric evaluation set based on the length of the shortest chain. We established a criterion where the shortest chain’s maximum length was set to 50 residues. Subsequently, we obtained a total of 61 instances of peptide-protein complexes.
@@ 2.2 Antigen-Antibody Version
We selected samples from the SAbDab [24] database with release dates between January 25, 2023, and August 9, 2023, as the evaluation set for antibody-related data, ensuring that the evaluation samples were not present in our training set. Further, the antigen type of peptide and protein were chosen. Additionally, we excluded samples containing more than 1400 residues. The test set comprised 131 antigen-antibody complexes and 13 nanobody-antigen complexes. The antibody chains from antigen-antibody complexes were extracted to serve as a test dataset for the evaluation of antibody variable regions (VH-VL). We extracted only the fragment variable regions of the heavy and light chains of the antibodies as inputs for the models.

### v2

@@ 1.4 Sequence Identity
To comprehensively grasp the impact of sequence similarity on model performance, we employ sequence identity as the similarity metric. Sequence identity is defined as the percentage of residues in the evaluation set chain that match those in the training set chain. We utilize the default configuration settings of the MMSeqs2[19] software suite for this computation.
@@ 2.1 General Version
Heteromeric protein complexes released from the period between January 16th, 2022, and December 12th, 2022, are gathered from Protein Data Bank (PDB) [25] as the evaluation set. Proteins with more than 1400 residues are excluded. These test samples are not present in the training set. The evaluation set for heteromeric proteins consists of 192 complexes.
The heteromeric protein test PDBs: 7a67 7aam 7acv 7acw 7acz 7al7 7b5g 7dz9 7dz9 7dz9 7dz9 7dz9 7dz9 7e1z 7e7c 7e9m 7eic 7eid 7eqc 7etr 7ewm 7f03 7f03 7f28 7f4n 7f4p 7f4s 7f4t 7fad 7fax 7fb8 7fba 7fc0 7fc0 7fc0 7fc0 7fc0 7fgj 7fiw 7kqk 7kql 7lo7 7mhr 7mrz 7n05 7n0a 7n0e 7nf8 7o0z 7ocj 7ocy 7od9 7oix 7oju 7oly 7ow1 7oxn 7pb4 7pb8 7pb8 7pb8 7pp2 7pqq 7pyt 7q4q 7q51 7qbe 7qbs 7qih 7qii 7qla 7qqy 7qu1 7quc 7quu 7qy5 7qy5 7r1z 7r1z 7r5a 7rc6 7rlw 7rlx 7rlz 7rm1 7rm3 7rqq 7rqr 7rsw 7ru6 7rx0 7rxq 7sch 7sd2 7sej 7sem 7sjp 7sk6 7sp8 7sp8 7srs 7ssc 7syy 7syy 7t2p 7t92 7t92 7t92 7tcq 7tcq 7tcu 7te1 7tgg 7tj4 7tr3 7u0a 7u7n 7u7n 7u8g 7ubz 7ulo 7unb 7unb 7unz 7urd 7urf 7uwj 7uym 7v4w 7vad 7vad 7vbq 7vcr 7vd7 7vgg 7vgr 7vkb 7vlz 7vmc 7wek 7wlw 7wmv 7wq3 7wq4 7wrs 7wuw 7wwq 7x2e 7x3c 7x4n 7x5c 7xjl 7y1r 7y3j 7y3j 7y6b 7y6c 7y6c 7y7m 7y8u 7ycu 7ypx 7yrh 7yx8 7yz9 7yzi 7z01 7zcl 7zk1 7zkw 7zkz 7zvi 7zvy 7zvy 7zw1 8a1e 8ap8 8ap8 8ap8 8b0u 8b38 8b38 8b38 8bd1 8d3v 8des 8dgq 8djk 8djk 8ds5 8gqe 8guo
To assess the models on protein-peptide complexes, we derive a subset of protein complexes from the heteromeric evaluation set based on the length of the shortest chain. We establish a criterion where the shortest chain’s maximum length is set to 50 residues. Subsequently, we obtain a total of 61 instances of peptide-protein complexes.
The protein-peptide test PDBs: 7aam 7acv 7acw 7e7c 7e9m 7eic 7eid 7fax 7fb8 7fba 7fc0 7fc0 7fc0 7fgj 7kqk 7n05 7oju 7ow1 7oxn 7pb4 7q4q 7q51 7qbs 7qqy 7quu 7qy5 7rc6 7rlw 7rlx 7rlz 7rm1 7rm3 7rqq 7rqr 7rsw 7rxq 7sjp 7ssc 7tcq 7u0a 7ulo 7urd 7urf 7uym 7v4w 7vcr 7vd7 7vlz 7wek 7wq3 7wq4 7wwq 7x2e 7x5c 7xjl 7y3j 7y3j 7y6c 7yx8 8des 8dgq
@@ 2.2 Antigen-Antibody Version
We select samples from the SAbDab [26] database with release dates between January 25, 2023, and August 9, 2023, as the evaluation set for antibody-related data, ensuring that the evaluation samples are not present in our training set. Further, the antigen type of peptide and protein are chosen. Additionally, we exclud samples containing more than 1400 residues. In cases where AlphaFold 3 predictions failed, we intersect the successful predictions from both AlphaFold 3 and HelixFold-Multimer. The test set comprises 141 antigen-antibody complexes. The antibody chains from antigen-antibody complexes are extracted to serve as a test dataset for the evaluation of antibody variable regions (VH-VL). We extract only the fragment variable regions of the heavy and light chains of the antibodies as inputs for the models.
The antibody-antigen test PDBs: 7u9e 8ium 8iw9 8dzv 8dy1 8iuk 8dy5 8axh 7yue 7ua2 8cz8 8db4 8hc4 8epa 8hc5 8hhy 8hhx 7uxl 7xdk 7xcz 7xda 8ek1 8eka 7xdl 8c3v 8bcz 8d7e 7wnb 7wn2 8i5i 8saw 7umn 7xj6 8dto 8sb0 8sb1 8saq 8sar 7zjl 8say 8i5h 8sb5 8sas 8sb3 8sb2 7uja 7uow 8ct6 8sav 7xj8 7xj9 8sax 7xik 8heb 7xs8 8hec 7xil 8hed 8a96 8gs9 8f0h 8fax 8fg0 8a99 7xsa 7xrz 8bse 7yru 8bsf 7zqt 8dn6 8hn6 8hn7 7st5 8dn7 8f6o 8gb8 8gb6 8ahn 8f60 8cwi 8cwj 8gb7 8dnn 8cwk 7xsc 8f6l 8av9 8g3n 8g3v 8g3r 7xeg 7xjf 8gnk 8g3q 8ee1 8ee0 8g3p 8g3o 8dwy 8g3z 8g3m 8dww 7quh 8g30 8h07 8czz 7xei 8a44 8d9y 8eoo 8da0 8da1 8d9z 8e1m 8scx 8smt 7zoz 8e1g 8duz 8elo 8elq 8de4 8cim 8elp 8gtp 7yd1 8byu 8gtq 7y8j 8el2 8ol9 7yds 8ib1 7yv1 8dz3 8dyx 7trh 8e6j 7yk4 8e6k